OpenStudy (anonymous):

Simplify the following rational expression completely. 36ab^2 / 72a^2bc

OpenStudy (anonymous):

\[\frac{ 36ab^2 }{ 72a^2bc }=\frac{ b }{ 2ac }\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

:) np